OpenStudy (anonymous):

For the sequence given by a(little 1)=3 and a(little n)=2a (little n-1) -4, find the first three terms

OpenStudy (anonymous):

@jim_thompson5910 @ajspeller

jimthompson5910 (jim_thompson5910):

what do you have so far

OpenStudy (anonymous):

Nothing, I honestly don't remember how to solve these

jimthompson5910 (jim_thompson5910):

What is the first term \(\Large a_{1}\) ?

OpenStudy (anonymous):

3?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

how about \(\Large a_{2}\) ?

OpenStudy (anonymous):

Uh...

OpenStudy (anonymous):

would it be 6??

jimthompson5910 (jim_thompson5910):

If you aren't sure, then replace EVERY copy of 'n' with 2 in \[\Large a_{n} = 2a_{n-1} - 4\]

jimthompson5910 (jim_thompson5910):

\[\Large a_{n} = 2a_{n-1} - 4\] \[\Large a_{{\color{red}{n}}} = 2a_{{\color{red}{n}}-1} - 4\] \[\Large a_{{\color{red}{2}}} = 2a_{{\color{red}{2}}-1} - 4\] \[\Large a_{2} = 2a_{1} - 4\]

OpenStudy (anonymous):

But how do you solve from that?

jimthompson5910 (jim_thompson5910):

well before you said a1 = 3, which is true so, \[\Large a_{2} = 2a_{1} - 4\] \[\Large a_{2} = 2(3) - 4\] \[\Large a_{2} = ??\]

OpenStudy (anonymous):

2?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

for a3, we do the same but now replace n with 3 \[\Large a_{n} = 2a_{n-1} - 4\] \[\Large a_{{\color{red}{n}}} = 2a_{{\color{red}{n}}-1} - 4\] \[\Large a_{{\color{red}{3}}} = 2a_{{\color{red}{3}}-1} - 4\] \[\Large a_{3} = 2a_{2} - 4\]

jimthompson5910 (jim_thompson5910):

in general, the idea is this (nth term) = 2*(previous term just before nth term) - 4

jimthompson5910 (jim_thompson5910):

so if you want the 10th term, you need to find the terms a1 through a9 since a10 relies on a9

OpenStudy (anonymous):

so to find 3 i put in 2 for a(little 2)??

jimthompson5910 (jim_thompson5910):

to find that third term, you need to use a2, yes

jimthompson5910 (jim_thompson5910):

as shown in \[\Large a_{3} = 2a_{2} - 4\]

OpenStudy (anonymous):

so 0 right?

jimthompson5910 (jim_thompson5910):

good, a3 = 0

jimthompson5910 (jim_thompson5910):

a1 = 3 a2 = 2 a3 = 0

OpenStudy (anonymous):

awesome!! thanks!

jimthompson5910 (jim_thompson5910):

np